OpenStudy (anonymous):

Find the work done by the force field F that is not conservative. The distance is between two points. How would you approach this problem since the Fundamental Rules of Calculus do no apply.

OpenStudy (anonymous):

integrate along the path F.dr

OpenStudy (anonymous):

Don't you need to Integrate?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

vector integral

OpenStudy (anonymous):

vector is <z,x,y> how would you do it when the path is a straight line?

OpenStudy (anonymous):

give us the question..well see

OpenStudy (anonymous):

between point (3,0,0) to point (0,pi/2,3)

OpenStudy (anonymous):

when the path between those two points is a straight line

OpenStudy (anonymous):

I'm having trouble knowing what to do when it's path dependent.

OpenStudy (anonymous):

whats the force?

OpenStudy (anonymous):

F(x,y,z)= z i + x j + y k

OpenStudy (anonymous):

gotcha

OpenStudy (anonymous):

so dr is the distance between the two points?

OpenStudy (anonymous):

first write the eqn of a line in cartesian form like x=3+3k y=k(pi)/2 z=3k

OpenStudy (anonymous):

parametric representation of the line segment right?

OpenStudy (anonymous):

the work done is F.dr which is <zdx, xdy, ydz> right

OpenStudy (anonymous):

sorry write y=-k(pi)/2 and z = -3k

OpenStudy (anonymous):

got it till here??

OpenStudy (anonymous):

Kinda it's just hard to remember how to find the parametric equations of a line

OpenStudy (anonymous):

no bt ive written em down fr u nw

OpenStudy (anonymous):

kk so use dot product of the vector with dr right?

OpenStudy (anonymous):

so you just replace x y z with the parametric equations right?

OpenStudy (anonymous):

so here goes dot prod gives zdx + xdy + ydz for zdx write (-3k)(dx) (-3k)(3dk)

OpenStudy (anonymous):

now integrate this -9k^2 dk from k =0 to k=-1

OpenStudy (anonymous):

coz k=-1 gives u ur final point

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

now similarly compute it for xdy and ydz and then add all three/

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

Yes so my area of integration is going to change because I reperamatize right?

OpenStudy (anonymous):

yes integrate all three frm 0 to -1

OpenStudy (anonymous):

bcoz remmbr on ur line k=0 gives the initial pt and k=-1 givs d final pt...so as our integrating variable is k, we use the lim its fr k

OpenStudy (anonymous):

that should do it i spose

OpenStudy (anonymous):

yup my book shows it but in two dimensions....

OpenStudy (anonymous):

js get the answer and tally it..its complicated enough...ur book mightve represented it in 2d..fr me dis is d way i thot dis cld be done

OpenStudy (anonymous):

wt hapnd??

OpenStudy (anonymous):

nothing I got the answer